# Bipartite Fluctuations of Critical Fermi Surfaces

### v2

@@ IV.1 Critical Theories
[…]
@@
For completeness, let us briefly mention the critical theory discussed in Ref. [55, 56], and illustrate its relation to the vortex theory Eq. (19) together with Eq. (26). The starting point is the parton construction Eq. (17) described in Sec. III. The bosonic parton $b$ is assumed to be further fractionalized into two fermions
@@
introducing an additional $\textrm{U}(1)$ gauge field denoted by $\widehat{a}$. One can assign the charge of $A-a$ to $f_{2}$, and put it in a mean-field state with Chern number $\mathsf{C}_{2}=1$. Then a Chern number changing phase transition of $f_{1}$ from $\mathsf{C}_{1}=-1$ to $\mathsf{C}_{1}=1$ will drive the transition of the chargon $b$ from a superfluid state to the bosonic Laughlin state at $\nu=1/2$. The band touching of $f_{1}$ typically involves two massless Dirac fermions $\chi_{1},\chi_{2}$. The full critical theory can be written as
[…]
@@
As we show in App. E, the critical theories described by Eq. (29) and Eq. (19) (together with Eq. (26)) can be related through fermionic particle-vortex duality [82, 83], where $\psi_{1},\psi_{2}$ can be interpreted as the fermionic vortices of $\chi_{1},\chi_{2}$. Consequently, the transition from the superfluid to the Laughlin state in the chargon sector can be effectively described by the integer QH transition of vortices. The analysis presented in Ref. [56] can be carried over for Eq. (26), demonstrating that other fermion bilinears are not permitted by translation symmetry under filling constraints.
@@ IV.2 Charge Fluctuations
Our primary focus is on understanding the static structure factor Eq. (8) at long wavelengths, as it governs the leading-order scaling of bipartite fluctuations $\mathcal{F}$ (see Tab. 1 and App. D). The subleading corner contribution in Eq. (12) or Eq. (16) depends on whether the localization tensor $\mathsf{G}^{ij}$, as described in Eq. (9) and Eq. (10), is finite or logarithmically divergent. In Landau FLs, it is well understood that the leading-order scaling follows $\mathcal{F}\sim L\log(L)$, while the subleading corner term is not well-defined due to the slow decay of spatial charge correlations [21]. This non-local feature of Landau FLs is tied to the power-law divergence of $\mathsf{G}^{ij}$. In the following, we will begin our analysis with the CFL phase and then proceed to examine the critical point.
@@ IV.2.1 CFL Phase
$\Pi_{f}^{\tau\tau}(k)$ $=-\mathscr{D}_{F}\left(1-\frac{|\omega|}{\sqrt{\omega^{2}+(v_{F}\bm{k})^{2}}}\right),$
$\Pi_{f}^{TT}(k)$ $=\mathscr{D}_{F}\frac{|\omega|\sqrt{\omega^{2}+(v_{F}\bm{k})^{2}}-\omega^{2}}{|\bm{k}|^{2}}.$ (30)
Here, $v_{F}$ is the fermi velocity, $k_{F}$ denotes the fermi momentum, and $\mathscr{D}_{F}=\frac{k_{F}}{2\pi v_{F}}$ represents the density of states at the fermi level. The vortex sector is in the integer QH state at the filling $\nu=-2$, which has the response function $\Pi_{v}(k)=-2\Pi_{\textrm{CS}}(k)=-2\frac{|\bm{k}|}{2\pi}\sigma^{1}$. The electron density-density correlation is determined using the Ioffe-Larkin rule Eq. (23). Upon integrating over all frequencies $\omega$, the static structure factor is obtained as follows
@@
$\approx\frac{1-e^{-(|\bm{k}|\ell_{B})^{2}/2}}{4\pi\ell_{B}^{2}}-\frac{|\bm{k}|^{3}\log(1/|\bm{k}|)}{4\pi^{2}k_{F}}+\ldots$ (31)
@@
The first term in Eq. (31) arises from the inter-Landau-level gapped modes [84, 85]. Based on calculations similar to those in Ref. [21], we find a corner contribution of $-\textrm{f}(\theta)/(8\pi^{2})$ to the bipartite charge fluctuations, where $\textrm{f}(\theta)$ is the universal angle function defined in Eq. (6). The second term $|\bm{k}|^{3}\log|\bm{k}|$ arises from the gapless modes in the lowest Landau level. This long-wavelength result has been predicted by both the HLR theory [63, 86] as well as the Son-Dirac theory [82, 87]. As we illustrate in App. D, the gapless modes contribute only to a boundary-law term. By combining the contributions from both gapped and gapless modes, we obtain the final result in Eq. (7) for the configuration depicted in Fig. 1. It is noteworthy that Eq. (7) applies to both the mentioned examples of gapped QH insulators [21] and gapless CFLs.
@@
It would be highly compelling if Eq. (7) holds true for all charge insulators, i.e., if the corner contribution is universally determined by the DC Hall conductivity $\sigma_{xy}$. However, unlike at the critical point, deep inside the insulator phase, the gapped modes exhibit short-range correlations and may be sensitive to microscopic details. As we clarify in a follow-up work [75], for systems with continuous translational and rotational symmetries, Eq. (12) generally holds, but with a non-universal value of the localization tensor $\mathsf{G}\geq|\sigma_{xy}|/2$. In other words, the coefficient of the universal angle function satisfies a universal lower bound set by the topological invariant $\sigma_{xy}$. In Ref. [75], we explain this bound and its saturation from the perspective of many-body quantum geometry [^4] Here, $\mathsf{G}^{ij}=\mathsf{G}\delta^{ij}$ serves as the many-body quantum metric [75], defined by adiabatic flux insertion (or twisted boundary conditions), where $i=x,y$ denotes the spatial indices.. It turns out that $\mathsf{G}=|\sigma_{xy}|/2$ is guaranteed under specific conditions, such as Galilean invariance and wavefunction holomorphicity, which are indeed satisfied by many examples. In fact, our predicted formula Eq. (7) has already been numerically verified by Monte Carlo simulations in Ref. [88] using the Rezayi-Read wavefunction, considering CFLs for fermions at $\nu=1/2$ and $\nu=1/4$, as well as CFLs for bosons at $\nu=1$ and $\nu=1/3$ fillings.
@@
More generally, the system may lack continuous translational and rotational symmetries, as in the models with spatially modulated periodic potentials discussed in Ref. [57, 58, 71], which concern the half-filled Chern band in Morié materials. In such cases, the corner contribution may depend on the corner angle, subregion orientation, and other microscopic details. Nevertheless, in Ref. [75], we demonstrate that the corner contribution can be expressed as an infinite series of orientation-resolved universal angle functions, generalizing the formula Eq. (6), along with their non-universal coefficients. However, these details are beyond the scope of this work.
@@ IV.2.2 Critical Point
Next, we turn our attention to the critical point, where the analysis becomes slightly more intricate. The QH transition in the vortex sector (i.e., the chargon sector) is presumably described by a CFT. It was argued in Ref. [55, 56] that the CFT sector and the $f$-fermion sector are dynamically decoupled right at the critical point. This assertion is supported by two key observations. (1) The Landau damping term of the gauge field $a_{\mu}$ effectively suppresses gauge fluctuations at the critical point. (2) When considering an operator $O$ from the CFT sector coupled to the $f$-fermion via $Of^{\dagger}f$, an additional Landau damping term $\int_{k}\frac{|\omega|}{|\bm{k}|}|O(k)|$ is generated. However, this term becomes irrelevant if the scaling dimension satisfies $\Delta[O]>3/2$, a condition likely met in the critical theory under consideration [55, 56].
Assuming the dynamical decoupling of the two sectors, the critical response in the vortex/chargon sector still maintains the CFT form [^5] In terms of space-time components, the response function has the form $\Pi_{\textrm{CFT}}^{\mu\nu}(k)=-\sigma_{xx}|k|(\delta_{\mu\nu}-\frac{k_{\mu}k_{\nu}}{|k|^{2}})-\sigma_{xy}\varepsilon_{\mu\nu\rho}k_{\rho}$.. Under the Coulomb gauge, it is expressed as
$\Pi_{\textrm{CFT}}(k)=\left(\begin{array}[]{cc}-\frac{|\bm{k}|^{2}}{\omega}\mathsf{F}_{L}(\frac{\omega}{|\bm{k}|})&-\sigma_{xy}|\bm{k}|\\ -\sigma_{xy}|\bm{k}|&\omega\mathsf{F}_{T}(\frac{\omega}{|\bm{k}|})\end{array}\right),$ (34)
@@
According to the Ioffe-Larkin composition rule Eq. (23) (or Eq. (80)), the static structure factor of electrons can be calculated as follows
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})=\int\frac{\textrm{d}\omega}{2\pi}\frac{|\bm{k}|^{2}\Pi_{f}^{\tau\tau}(k)(\mathsf{F}_{L}^{b}(\frac{\omega}{|\bm{k}|})(\omega\mathsf{F}_{T}^{b}(\frac{\omega}{|\bm{k}|})+\Pi_{f}^{TT}(k))-\omega(\sigma_{xy}^{b})^{2})}{(\omega\mathsf{F}_{T}^{b}(\frac{\omega}{|\bm{k}|})+\Pi_{f}^{TT}(k))(|\bm{k}|^{2}\mathsf{F}_{L}^{b}(\frac{\omega}{|\bm{k}|})-\omega\Pi_{f}^{\tau\tau}(k))-\omega|\bm{k}|^{2}(\sigma_{xy}^{b})^{2}}=-\frac{\sigma_{xx}^{b}}{\pi}|\bm{k}|^{2}\log(1/|\bm{k}|)+\ldots$ (37)
@@
Here, $\Pi_{f}^{\tau\tau}(k)$ and $\Pi_{f}^{TT}(k)$ are the components of the response function of $f$-fermions given in Eq. (30). The scaling functions $\mathsf{F}_{L}^{b}(k)$ and $\mathsf{F}_{T}^{b}(k)$ are provided by Eq. (35), with the coefficient $\sigma_{xx}^{b}$. Evaluating the frequency integral in Eq. (37) is not an easy task. In App. F, we present two methods for extracting the long-wavelength behavior $|\bm{k}|^{2}\log(1/|\bm{k}|)$. By retaining only the leading-order term of the integrand in the small-$\bm{k}$ expansion, the $\omega$-integral can be performed analytically. Additionally, we have numerically evaluated the full expression in Eq. (37). Remarkably, at small $|\bm{k}|/k_{F}$, the numerical result aligns closely with the analytical expression.
@@
The Fourier transform of Eq. (37) yields the power-law spatial correlation given by Eq. (2), with an exponent $\alpha=4$. The overall coefficient $C_{0}$ is given by the universal constant $C_{\rho}=2(\sigma_{xx}^{b})/\pi^{2}$, which identifies the current central charge of the CFT describing the chargon sector. By examining the integrand in Eq. (37) and sequentially taking the limits $\omega\rightarrow 0$ and then $\bm{k}\rightarrow 0$, we can demonstrate the vanishing compressibility. Moreover, the vanishing Drude weight can be confirmed using the Ioffe-Larkin rule Eq. (23) (or Eq. (80)). As a result, we identify an incompressible state with vanishing Drude weight, characterized by a logarithmically divergent localization tensor as described by Eq. (14), despite the absence of conformal symmetry.
@@
According to Refs. [20, 18, 19, 21] (also see App. D), for the subregion depicted in Fig. 1, the bipartite charge fluctuations follow Eq. (16). Similar to the first term in Eq. (31), note that Eq. (37) also includes a term $|\bm{k}|^{2}$ arising from gapped modes, meaning that the localization tensor contains subleading contributions in addition to the divergence described in Eq. (14). These modes contribute to a corner contribution $(\textrm{const})\textrm{f}(\theta)$ that does not scale with $L$, similar to the subleading term in Eq. (12). The coefficient of this term is a function of $\sigma_{xx}^{b}$ and $\sigma_{xy}^{b}$. However, isolating this subleading constant term from the $\log(L)$ scaling in experiments or numerical simulations can be challenging. Additionally, in theory, this constant term is also sensitive to the UV cutoff.
@@
One may wonder whether the universal number $C_{\rho}$ can be detected through other experimental observables, such as transport measurements. A promising candidate is the predicted resistivity jump at the critical point [56], which is solely determined by the universal data of the CFT sector. The key argument leading to the universal resistivity jump is based on the observation that $\Pi^{\tau j}(\mathtt{i}\omega,\bm{k}\rightarrow 0)=0$, where $j=x,y$, holds for both the $f$-fermion sector and the $b$-chargon sector. Considering the spatial components of the Ioffe-Larkin rule Eq. (23) (or Eq. (80)), the total resistivity is given by $\rho_{ij}=\rho^{f}_{ij}+\rho^{b}_{ij}$. As one approaches the critical point from the Landau FL phase, additional contributions to the longitudinal and Hall components arise from the gapless degrees of freedom in the CFT:
@@
These equations allow us to solve for $\sigma_{xx}^{b}$ and express the universal coefficient $C_{\rho}=2(\sigma_{xx}^{b})/\pi^{2}$ in Eq. (16) as a function of $\Delta\rho_{xx}$ and $\Delta\rho_{xy}$, as in Eq. (15). Note that both $\Delta\rho_{xx}$ and $\Delta\rho_{xy}$ are scaling functions of $\omega/T$ (i.e., frequency over temperature), and Eq. (15) should be understood as a relation in the limit $T\rightarrow 0$. The values of $\Delta\rho_{xx}$ and $\Delta\rho_{xy}$, obtained through the large-$N$ expansion, are available in Refs. [89, 76], and could potentially be determined more precisely using the numerical method of conformal bootstrap [64].

### v3

@@ IV.1 Critical Theories
$\mathcal{L}_{\textrm{vrtx}}[\psi,\tilde{a}]=\sum_{I=1}^{2}\bar{\psi}_{I}\not{\textrm{D}}_{\tilde{a}}\psi_{I}+\frac{1}{2e^{2}}\textrm{d}\tilde{a}\wedge\star\textrm{d}\tilde{a}$, (29)
@@
For completeness, let us briefly mention the critical theory discussed in Refs. [56, 57], and illustrate its relation to the vortex theory Eq. (19) together with Eq. (29). The starting point is the parton construction Eq. (17) described in Sec. III. The bosonic parton $b$ is assumed to be further fractionalized into two fermions
@@
introducing an additional $\textrm{U}(1)$ gauge field denoted by $\widehat{a}$. One can assign the charge of AaA-a to $f_{2}$, and put it in a mean-field state with Chern number $\mathsf{C}_{2}=1$. Then a Chern number changing phase transition of $f_{1}$ from $\mathsf{C}_{1}=-1$ to $\mathsf{C}_{1}=1$ will drive the transition of the chargon $b$ from a superfluid state to the bosonic Laughlin state at $\nu=1/2$. The band touching of $f_{1}$ typically involves two massless Dirac fermions $\chi_{1},\chi_{2}$. The full critical theory can be written as
$\mathcal{L}$ $=\mathcal{L}_{\textrm{FS}}[f,a]+\sum_{I=1}^{2}\bar{\chi}_{I}\not{\textrm{D}}_{\widehat{a}}\chi_{I}-\frac{2\mathtt{i}}{4\pi}\widehat{a}\wedge\textrm{d}\widehat{a}$
@@
As we show in App. F, the critical theories described by Eq. (32) and Eq. (19) (together with Eq. (29)) can be related through fermionic particle-vortex duality [94, 95], where $\psi_{1},\psi_{2}$ can be interpreted as the fermionic vortices of $\chi_{1},\chi_{2}$. Consequently, the transition from the superfluid to the Laughlin state in the chargon sector can be effectively described by the integer QH transition of vortices. The analysis presented in Ref. [57] can be carried over for Eq. (29), demonstrating that other fermion bilinears are not permitted by translation symmetry under filling constraints.
@@ IV.2 Charge Fluctuations
Our primary focus is to understand the static structure factor Eq. (8) at long wavelengths, which governs the leading-order scaling of bipartite fluctuations $\mathcal{F}$ (see Tab. 1 and App. D). Whether a subleading corner contribution appears, as in Eq. (12) or Eq. (16), depends on whether the localization tensor $\mathsf{G}^{ij}$, as defined in Eq. (9) and Eq. (10), is finite or logarithmically divergent. In Landau FLs, it is well understood that the leading-order scaling follows $\mathcal{F}\sim L\log(L)$, while the subleading corner contribution is not well-defined due to the slow decay of spatial charge correlations [22]. This non-local feature of Landau FLs is directly tied to the power-law divergence of $\mathsf{G}^{ij}$. In the following, we first analyze the CFL phase and then proceed to examine the critical point.
@@ IV.2.1 CFL Phase
$\Pi_{f}^{\tau\tau}(k)$ $=\mathscr{D}_{F}\left(1-\frac{|\omega|}{\sqrt{\omega^{2}+(v_{F}\bm{k})^{2}}}\right),$
$\Pi_{f}^{TT}(k)$ $=-\mathscr{D}_{F}\frac{|\omega|\sqrt{\omega^{2}+(v_{F}\bm{k})^{2}}-\omega^{2}}{|\bm{k}|^{2}}.$ (33)
Here, $v_{F}$ is the fermi velocity, $k_{F}$ denotes the fermi momentum, and $\mathscr{D}_{F}=\frac{k_{F}}{2\pi v_{F}}$ represents the density of states at the fermi level. The vortex sector is in the integer QH state at the filling $\nu=-2$, which has the response function $\Pi_{v}(k)=-2\Pi_{\textrm{CS}}(k)=2\frac{|\bm{k}|}{2\pi}\sigma^{1}$. The electron density-density correlation is determined using the Ioffe-Larkin rule Eq. (23). Upon integrating over all frequencies $\omega$, the static structure factor is obtained as follows
@@
$\approx\frac{1-e^{-(|\bm{k}|\ell_{B})^{2}/2}}{4\pi\ell_{B}^{2}}+\frac{|\bm{k}|^{3}\log(1/|\bm{k}|)}{4\pi^{2}k_{F}}+\ldots$ (34)
@@
The first term in Eq. (34) originates from the inter-Landau-level gapped modes [96, 97]. Following a calculation similar to that in Ref. [22], we find that these modes contribute a corner term of $\textrm{f}(\theta)/(8\pi^{2})$ to the bipartite charge fluctuations, where $\textrm{f}(\theta)$ is the universal angle function defined in Eq. (6). The second term, proportional to $|\bm{k}|^{3}\log(1/|\bm{k}|)$, arises from the gapless modes in the lowest Landau level. This long-wavelength result has been predicted by both the HLR theory [67, 98] as well as the Son-Dirac theory [94, 99]. As shown in App. D, the gapless sector contributes only to a boundary-law term. By combining the effects of both gapped and gapless modes, we arrive at the final expression in Eq. (7) for the configuration shown in Fig. 1. Notably, Eq. (7) holds for both gapped QH insulators [22] and gapless CFLs.
@@
It would be highly compelling if Eq. (7) holds true for all charge insulators, i.e., if the corner contribution is universally determined by the DC Hall conductivity $\sigma_{xy}$. However, unlike at the critical point, deep inside the insulator phase, the gapped modes exhibit short-range correlations and may be sensitive to microscopic details. As we clarify in a follow-up work [71], for systems with continuous translational and rotational symmetries, Eq. (12) generally holds, but with a non-universal value of the localization tensor $\mathsf{G}\geq|\sigma_{xy}|/2$. In other words, the coefficient of the universal angle function satisfies a universal lower bound set by $\sigma_{xy}$. In Ref. [71], we explain this bound and its saturation from the perspective of many-body quantum geometry [^5] For gapped insulators, $\mathsf{G}^{ij}=\mathsf{G}\delta^{ij}$ serves as the many-body quantum metric [71], defined by adiabatic flux insertion (or twisted boundary conditions), where $i=x,y$ denotes the spatial indices. (see also App. E). It turns out that $\mathsf{G}=|\sigma_{xy}|/2$ is guaranteed under specific conditions, such as Galilean invariance and wavefunction holomorphicity, which are indeed satisfied by many examples. In fact, our predicted formula Eq. (7) has already been numerically verified by Monte Carlo simulations in Ref. [100] using the Rezayi-Read wavefunction, considering CFLs for fermions at $\nu=1/2$ and $\nu=1/4$, as well as CFLs for bosons at $\nu=1$ and $\nu=1/3$ fillings.
@@
More generally, the system may lack continuous translational and rotational symmetries, as in the models with spatially modulated periodic potentials discussed in Refs. [58, 59, 81], which concern the half-filled Chern band in morié materials. In such cases, the corner contribution may depend on the corner angle, subregion orientation, and other microscopic details. Nevertheless, in Ref. [71], we demonstrate that the corner contribution can be expressed as an infinite series of orientation-resolved universal angle functions, generalizing the formula Eq. (6), along with their non-universal coefficients. However, these details are beyond the scope of this work.
@@ IV.2.2 Critical Point
Next, we turn our attention to the critical point, where the analysis becomes slightly more involved. The QH transition in the vortex sector (i.e., the chargon sector) is presumably described by a CFT. As argued in Refs. [56, 57], the CFT sector and the $f$-fermion sector are dynamically decoupled precisely at the critical point—a reasoning we have reviewed in Sec. III.2.
As a result, the critical response in the vortex (or chargon) sector retains its CFT form [^6] In terms of space-time components, the current-current correlation has the form $\Pi_{\textrm{CFT}}^{\mu\nu}(k)=\sigma_{xx}|k|(\delta_{\mu\nu}-\frac{k_{\mu}k_{\nu}}{|k|^{2}})+\sigma_{xy}\varepsilon_{\mu\nu\rho}k_{\rho}$.. Under the Coulomb gauge, it is expressed as
$\Pi_{\textrm{CFT}}(k)=\left(\begin{array}[]{cc}\frac{|\bm{k}|^{2}}{\omega}\mathsf{F}_{L}(\frac{\omega}{|\bm{k}|})&\sigma_{xy}|\bm{k}|\\ \sigma_{xy}|\bm{k}|&-\omega\mathsf{F}_{T}(\frac{\omega}{|\bm{k}|})\end{array}\right),$ (37)
@@
According to the Ioffe-Larkin rule Eq. (88) or Eq. (23), the static structure factor can be calculated as follows
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})=\int\frac{\textrm{d}\omega}{2\pi}\frac{|\bm{k}|^{2}\Pi_{f}^{\tau\tau}(k)(\mathsf{F}_{L}^{b}(\frac{\omega}{|\bm{k}|})(\Pi_{f}^{TT}(k)-\omega\mathsf{F}_{T}^{b}(\frac{\omega}{|\bm{k}|}))+\omega(\sigma_{xy}^{b})^{2})}{(\Pi_{f}^{TT}(k)-\omega\mathsf{F}_{T}^{b}(\frac{\omega}{|\bm{k}|}))(|\bm{k}|^{2}\mathsf{F}_{L}^{b}(\frac{\omega}{|\bm{k}|})+\omega\Pi_{f}^{\tau\tau}(k))+\omega|\bm{k}|^{2}(\sigma_{xy}^{b})^{2}}=\frac{\sigma_{xx}^{b}}{\pi}|\bm{k}|^{2}\log(1/|\bm{k}|)+\ldots$ (40)
@@
Here, $\Pi_{f}^{\tau\tau}(k)$ and $\Pi_{f}^{TT}(k)$ are the components of the response function of $f$-fermions given in Eq. (33). The scaling functions $\mathsf{F}_{L}^{b}(k)$ and $\mathsf{F}_{T}^{b}(k)$ are provided by Eq. (38), with the coefficient $\sigma_{xx}^{b}$. Evaluating the frequency integral in Eq. (40) is not an easy task. In App. G, we present two methods for extracting the long-wavelength behavior $|\bm{k}|^{2}\log(1/|\bm{k}|)$. By retaining only the leading-order term of the integrand in the small-$\bm{k}$ expansion, the $\omega$-integral can be performed analytically. Additionally, we have numerically evaluated the full expression in Eq. (40). Remarkably, at small $|\bm{k}|/k_{F}$, the numerical result aligns closely with the analytical expression.
@@
The Fourier transform of Eq. (40) yields the power-law spatial correlation given by Eq. (2), with an exponent $\alpha=4$. The overall coefficient $C_{0}$ is given by the universal constant $C_{\rho}=2(\sigma_{xx}^{b})/\pi^{2}$, which identifies the current central charge of the CFT describing the chargon sector. By examining the integrand in Eq. (40) and sequentially taking the limits $\omega\rightarrow 0$ and then $\bm{k}\rightarrow 0$, we can demonstrate the vanishing compressibility. Moreover, the vanishing Drude weight can be confirmed using the Ioffe-Larkin rule Eq. (88) or Eq. (23). As a result, we identify an incompressible state with vanishing Drude weight, characterized by a logarithmically divergent localization tensor as described by Eq. (14), despite the absence of conformal symmetry.
@@
According to Refs. [20, 18, 19, 22] (see also App. D), for the subregion depicted in Fig. 1, the bipartite charge fluctuations follow the CFT-like expression in Eq. (16). Similar to the first term in Eq. (34), Eq. (40) also includes a $|\bm{k}|^{2}$ contribution from gapped modes, implying that the localization tensor contains finite subleading terms in addition to the divergence in Eq. (14). These modes contribute to a corner term of the form $(\textrm{const})\textrm{f}(\theta)$ that does not scale with $L$, similar to the subleading contribution in Eq. (12). The coefficient of this term depends on $\sigma_{xx}^{b}$ and $\sigma_{xy}^{b}$. However, isolating this subleading constant from the $\log(L)$ scaling in experiments or numerical simulations can be challenging. Theoretically, this constant term is also sensitive to the UV cutoff.
@@
One may wonder whether the universal number $C_{\rho}$ can also be probed through other experimental observables, such as transport measurements. A promising candidate is the predicted resistivity jump at the critical point [57], which is entirely determined by universal data from the CFT sector. The key argument [57] leading to this universal resistivity jump is based on the observation that $\Pi^{\tau j}(\mathtt{i}\omega,\bm{k}\rightarrow 0)=0$, where $j=x,y$, holds for both the $f$-fermion and the $b$-chargon sectors. Considering the spatial components of the Ioffe-Larkin rule Eq. (88) or Eq. (23), the total resistivity is given by $\rho_{ij}=\rho^{f}_{ij}+\rho^{b}_{ij}$. As one approaches the critical point from the Landau FL phase, additional contributions to the longitudinal and Hall components arise from the gapless degrees of freedom in the CFT:
@@
These expressions allow one to solve for $\sigma_{xx}^{b}$ and express the universal coefficient $C_{\rho}=2(\sigma_{xx}^{b})/\pi^{2}$ in Eq. (16) as a function of $\Delta\rho_{xx}$ and $\Delta\rho_{xy}$, as shown in Eq. (15). Note that the resistivity tensor $\Delta\rho_{ij}$ is a scaling function of $\omega/T$ (i.e., frequency over temperature), and Eq. (15) should be interpreted as a relation in the limit $T\rightarrow 0$. The values of $\Delta\rho_{xx}$ and $\Delta\rho_{xy}$, computed via large-$N$ expansions, are available in Refs. [101, 86], and could potentially be determined more precisely using conformal bootstrap techniques [68, 69].
We conclude this section by emphasizing that the predicted universal behaviors in Eq. (14) and Eq. (16) are robust and remain independent of the specific form of the Fermi-surface response function $\Pi_{f}$. For illustration purposes, we have used the special case given by Eq. (33) in the integral Eq. (40). However, in general, $\Pi_{f}$ is determined by the electron band structure and may exhibit anisotropy and sensitivity to UV details. Despite this, Eq. (14) and Eq. (16) remain valid. This is because, for the purpose of evaluating corner charge fluctuations, one essentially has $\Pi^{-1}=\Pi_{f}^{-1}+\Pi_{b}^{-1}\approx\Pi_{b}^{-1}$, where $\Pi_{b}$ takes the CFT expression Eq. (37) due to the emergent conformal symmetry in the vortex/chargon sector at criticality. This result follows as a general consequence of the Ioffe-Larkin rule and the distinct scaling behaviors of $\Pi_{b}$ and $\Pi_{f}$ at the critical point.